OpenStudy (anonymous):

Simplify: (3x − 5) + (3x + 6) 6x − 1 1 6x − 11 6x + 1

OpenStudy (anonymous):

another "simplify" it means "add"\[(3x − 5) + (3x + 6)=3x+3x-5+6=6x+1\]

OpenStudy (anonymous):

(3x-5) + (3x+6) step 1: combine like terms (3x +3x)=6x+1=(6-5) so the answer is 6x+1

mathslover (mathslover):

You have : (3x-5) + (3x+6) First, open the brackets : \(\color{blue}{3x - 5} + \color{red}{3x+6}\) Now notice the two "3x" , combine the like terms : \(\color{blue}{3x+3x} \color{red}{ -5 + 6}\) Now, solve for the blue ones first and then for the red ones. 3x + 3x = ? - 5 + 6 = ?

mathslover (mathslover):

@SOS101 , what is 3x + 3x ?

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

@HelloGoodmorning Please give the medal to the person that helped me , they all did but whoever you want to choose

OpenStudy (anonymous):

Will do

OpenStudy (anonymous):

Decision had been made

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

@mathslover congratulations my friend congrats

OpenStudy (anonymous):

Your very welcome @sos101

OpenStudy (anonymous):

Don't forget to fan me everyone have a nice blessed life